3- Design, Implement a combinational circuit that generates the equation: Y=b2+4 Assuming input b as 3-bit binary number denoAssignment (2) Repeat Assignment (2) using 3X8 line decoder in the implementation Process
Answer #1

. here maximum 3 bit numby = 7(111) =) max ) = 49+4= 53 28-32 ; 26.64 Aso we need ) b y =b²tu (6): for y / ~ y Ys Yu Y Y Y Yo0-2- DY DY - ③ by wing dewday we can directly to it Ty is of if it is thisif you have any doubts you can comment below